OpenStudy (anonymous):

find lim x->0 ln(x+1)/x using the derivative

OpenStudy (ash2326):

we have \[\lim_{x\to 0}\frac{\ln (x+1)}{x}\] if we put x=0 then we get \[\frac 0 0\] so we can apply L'Hopitol here. Take derivative of both numerator and denominator, we get \[\huge \lim_{x\to 0}\frac{\frac 1{(x+1)}}{1}\] can you solve now?

OpenStudy (anonymous):

wait a minute i think i found the solution for this, recalling some stuff from pre-calculus

OpenStudy (ash2326):

*L'Hopital 's

OpenStudy (anonymous):

cant't use L'Hopsital rule

OpenStudy (anonymous):

i know i s 1

OpenStudy (ash2326):

you mentioned that it has to be done using the derivative

OpenStudy (anonymous):

correct

OpenStudy (ash2326):

I did the same thing

OpenStudy (anonymous):

but forget to mention that u can't use L'Hosp for this proble

OpenStudy (anonymous):

i recall that ln (m/n ) u can change to ln m- ln n

OpenStudy (anonymous):

i guess this problem is involve with chain rule

OpenStudy (ash2326):

I got something

OpenStudy (ash2326):

We know \[f'(x)=\lim_{h\to 0} \frac{f(x+h)-f(x)}{x+h-x}\]

OpenStudy (anonymous):

k that may work

OpenStudy (ash2326):

let \[f(x) =\ln (x+1)\] so \[f'(x)=\lim_{h\to 0} \frac{\ln (x+h+1)-\ln(x+1)}{x+h-x}\] we know that f'(x)= 1/(x+1 )so \[\frac{1}{x+1}=\lim_{h\to 0} \frac{\ln (x+h+1)-\ln(x+1)}{x+h-x}\] \[\large \frac{1}{x+1}=\lim_{h\to 0} \frac{\ln (\frac{x+h+1}{x+1})}{h}\] we get \[\large \frac{1}{x+1}=\lim_{h\to 0} \frac{\ln (\frac{h}{x+1}+1)}{h}\] put x=0 \[\large \frac{1}{1}=\underline{\lim_{h\to 0} \frac{\ln (\frac{h}{1}+1)}{h}}\] this is our limit so, we have proved that it's equal to 1 just replace h by x, cause it doesn't depend on variable

OpenStudy (ash2326):

@Andresfon12 do you understand this?

OpenStudy (anonymous):

yes @ash2326

OpenStudy (ash2326):

good :)

OpenStudy (zzr0ck3r):

very clever

mathslover (mathslover):

nice explanation @ash2326 ..